OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch apothem A = sq. in.

OpenStudy (mind2000):

what is the area of an equilateral triangle (regular 3-gon) with the given measurement? 6-inch apothem, apothem a = 6 an equilateral triangle has 6 30-60-90 triangles each with the shortest leg being the apothem, length = a = 6 find hypotenuse of each of the 6 triangles: angle opposite to a is 30 degrees, sin 30 = opposite/hypotenuse = a/hyp = 6/hyp 0.5 = 6/hyp 0.5hyp = 6 hyp = 12 inches find adjacent of each of the 6 triangles: cos 30 = adjacent/hypotenuse = adj/12 sqrt(3)/2 = adj/12 12sqrt(3)/2 = adj 6sqrt(3) = adj find side of the equilateral triangle: adj is 1/2 the length of a side of the equilateral triangle side s = 2 * adj = 2 * 6sqrt(3) = 12sqrt(3) find perimeter p of the equilateral triangle: p = 3 * s = 3 * 12sqrt(3) = 36sqrt(3) find area of one of the 6 30-60-90 triangles: area of one of the 6 triangles = 1/2 * adj * a area of one of the 6 triangles = 1/2 * 6sqrt(3) * 6 = 18sqrt(3) find area of equilateral triangle: area, A, of equilateral triangle = 6 * 18sqrt(3) A = 108sqrt(3) square inches A = 187.1 square inches rounded to nearest tenth find area equation for any equilateral triangle given apothem: A = 108sqrt(3) = 3 * 36sqrt(3) = 3 * perimeter of the equilateral triangle = a/2 * p A = 1/2 * a * p (p = perimeter, A = area, a = apothem, s = side) (p = 3 * s) A = 1/2 * a * 3 * s s = 12sqrt(3) = 6 * 2sqrt(3) = 2a * sqrt(3) A = 3/2 * a * 2a * sqrt(3) A = 3 * a^2 * sqrt(3)